BTC Inc. has the following financial statements:

$m 2018 2017
Current Assets $        245 $        210
Net Fixed Assets $          67 $          42
Total Assets $        312 $        252
Current Liabilities $          78 $          60
LT Debt $          76 $          64
Total Liabilities $        154 $        124
Common Equity & Ret. Earn $        158 $        128
Total Liab. And Equity $        312 $        252

Revenue in 2018 was $455m and EBITDA was 15% of revenue. Depreciation was 11% of Net Fixed Assets and Corp. Tax was 35%. Free Cash Flow (FCF) for 2018 was:

Select one:

a. $45m

b. $28m

c. None of the Above

d. $64m

e. $32m
Answer #1

FCF = (1-TAX)*(EBITDA+Depreciation-Change in Net working capital - capital expenditures

EBITDA = 15% of revenue = 15%*455 = 68.25

Depreciation = 11% of net fixed assets in 2018 = 11%*67 = 7.37

change in net working capital = net working capital in 2018 - net working capital in 2017

= CA-CL in 2018 - CA-CL in 2017

= (245-78)-(210-60) = 17

capital expenditure = change in net fixed assets

=67-42 = 25

FCF = (68.25+7.37-17-25)*(1-35%) = 21.853

aNS: None of the above